Subtract the sum of (4X^2 +3X-1) and (3X^2 -1) From (8X^2 -5X + 2)
Answer by uday1435(57)

Here we have to add (4x^2 + 3x -1 ) and (3x^2 -1)
We can add up only the coefficients of the like terms
Which means
(4x^2 + 3x -1 ) + (3x^2 -1) = (4+3) x^2 +3x (–1 – 1) = 7x^2 + 3x -2
We need to subtract this (7x^2 + 3x -2) from (8x^2 –5x +2)
Let us do that we need to change the signs (shown in red) of all terms in (7x^2 + 3x -2)
(8x^2 –5x +2) - (7x^2 + 3x -2) = 8^2 –5x +2 – 7x^2 – 3x+2
= (8- 7)x^2 (- 5 – 3)x (+2+2) = x^2 – 8x + 4